# DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models

### v1

@@ A.4 The Application Scope of Hallucination Labels
means there is a contradiction between the dialogue role and his context, a contradiction with the dialogue object, or an internal contradiction within the current response (in a declarative sentence). It also refers to factual and relevant nonsense or contextually inappropriate responses.
@@
involves responses that are irrelevant to the dialogue topic or are unrelated due to a misunderstanding (in the case of a question).
@@
is that the LLM excessively relies on the correctness of context, generating responses for unanswerable content (in a declarative sentence).
@@
Benchmark By LLMs Faith Halu Dialogue Explanation
FactCollect - - - -
HADES - $\checkmark$ - $\checkmark$
FactCHD $\checkmark$ - - $\checkmark$
HaluEval - - $\checkmark$ $\checkmark$
WikiBio+ $\checkmark$ - - -
PHD $\checkmark$ - - -
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 5: The comparison between our DiaHalu and other benchmarks. ‘By LLMs’, ‘Faith Halu’, ‘Dialogue’, and ‘Explanation’ mean whether it is naturally generated by LLMs, whether it provides faithfulness hallucination, whether it is at dialogue level, and whether there are explanations respectively.
A.5 The Supplementary Details for Annotation
Figure 11: The annotation interface on Label Studio.
A.6 The Prompt Designed for Detection
In Figure 12, we show the whole prompt specially designed for hallucination detection of the baselines: LLaMA-30B, Vicuna-33B, ChatGPT3.5, and ChatGPT4. It is worth noting that due to the poor instruction-following ability and the disorderly output format of the LLaMA and Vicuna models, we conduct experiments in a 1-shot manner. The ChatGPT3.5 used for detection is ChatGPT3.5 (gpt-3.5-turbo-1106) [^6] https://platform.openai.com/docs/guides/text-generation/chat-completions-api and the ChatGPT4 is ChatGPT4 (gpt-4-1106-preview)  [^7] https://openai.com/gpt-4.
@@ A.6 The Prompt Designed for Detection
A.7 The Settings for CoT and Retrieval
We employ CoT in all four domains to enhance the performance of ChatGPT3.5 and ChatGPT4. The specific CoT is illustrated in Figure 13. While only knowledge-grounded and reasoning domains are tested with retrieval via Google [^8] https://console.cloud.google.com/apis/library. This is because domains of task-oriented and chit-chat mainly involve scenarios related to daily life or virtual worlds, without specific domain knowledge as supplementary information.
@@ A.7 The Settings for CoT and Retrieval
A.8 Future Works
@@
is an important work in the future. We propose the first dialogue-level hallucination detection evaluation benchmark and experimental results show that it is a very challenging task. Therefore, developing methods based on this dataset to achieve a relatively high recognition accuracy is highly valuable.
@@
is an extension task of this work. Most existing hallucination elimination methods primarily focus on passage-level factuality hallucination (Luo et al., 2024). Hallucination elimination at the dialogue level not only requires the models to have a long-context memory capability but also demands the abilities to recognize changes in topics and logical transitions in the dialogue. These abilities are also helpful in addressing faithfulness hallucinations. At the same time, improving the accuracy of knowledge in the LLM parameters and the reasoning ability are equally important.
@@
is the phenomenon that LLMs tend to accumulate hallucinations rather than self-correcting during the generation process (Zhang et al., 2023). Some previous works also validate this phenomenon (Azaria and Mitchell, 2023; Ang et al., 2023). In our dataset, there is a noticeable issue of hallucination snowballing, so we only label the first occurrence of hallucination in our dataset. Eliminating hallucination snowballing is also important in the future.
@@
During the annotation process of this dataset, we introduce a hallucination category termed "overreliance", which represents answering unanswerable content (Slobodkin et al., 2023; Sulem et al., 2021). This phenomenon signifies that LLMs tend to trust the input provided by users. Sometimes, even when there are errors in user input, the LLMs still fail to recognize them. A few past researches explore the related areas. However, the issue in the application of human-machine interaction and multi-agent scenarios remains crucial.

### v2

@@ A.7 The Application Scope of Hallucination Labels
means there is a contradiction between one of the speakers and himself in the dialogue context, a contradiction between the two speakers (when both are declarative sentences) in the dialogue context. It also refers to factual and relevant nonsense, contextually inappropriate responses or other inconsistent errors.
@@
involves responses that are irrelevant to the dialogue topic or an irrelevance due to misunderstanding the grammar of a question. (Please note that we emphasize the use cases for both interrogative and declarative sentences.)
@@
is that the LLM excessively trust in the correctness of the context, generating serious responses to statements that were inherently wrong or unanswerable (in a declarative sentence).
@@
A.8 Baselines and Metrics
I. Baselines Selected
Below is a detailed description of all the baselines we selected.
Random
FaithCritic
(Dziri et al., 2022a) is one of the most effective dialog text hallucination classifiers before the era of large language models. Trained on a large-scale dialog corpus, it can output the confidence level for each classification label. Since the model’s input includes dialog-related knowledge, we use the retrieved contents as the knowledge during the experiment.
SelfCheckGPT
FOCUS
(Zhang et al., 2023b) is an improved version of SelfCheckGPT. It takes into account the attention scores between entity tokens, enabling more accurate classification of hallucination at both the sentence and paragraph levels.
LLaMa-30B && Vicuna-33B
They are two well-pretrained and widely deployed open-source LLM backbones [^5] https://huggingface.co/huggyllama/llama-30b, https://huggingface.co/lmsys/vicuna-33b-v1.3 (Touvron et al., 2023; Chiang et al., 2023). We provide a specially designed prompt to assist with detection. More details about this prompt are shown in Appendix A.9.
Gemini1.5 PRO
Gemini1.5 PRO [^6] https://gemini.google.com/ (Anil et al., 2023) is the latest version of the language model launched by Google. It inherits the powerful natural language processing capabilities of its predecessor and has made significant improvements in understanding and generating text. We employ the same prompt for binary detection as LLaMa-30B and Vicuna-33B do. We also create a manually prompt to assist with fine-grained recognition in Appendix A.10.
ChatGPT3.5 && GPT4
Both of these models are developed by OpenAI [^7] https://chat.openai.com/. ChatGPT3.5 marks the beginning of the era of large language models and GPT4 is currently the most powerful language model (Wu et al., 2023c; OpenAI, 2023). We employ the same prompt as LLaMa-30B and Vicuna-33B do for binary detection. And the same prompt as Gemeni1.5 PRO do is used for fine-grained recognition. The ChatGPT3.5 version is ChatGPT3.5 (gpt-3.5-turbo) [^8] https://platform.openai.com/docs/models/gpt-3-5-turbo and the GPT4 version is GPT4 (gpt-4-turbo)  [^9] https://openai.com/gpt-4.
II. Metrics Calculation
Despite we annotating the subtypes of hallucination in the dataset, achieving consistent labels even among humans requires further discussion. Therefore, similar to past hallucination detection efforts, we first focus on a binary classification task of determining the existence of hallucination. Consequently, we utilize binary classification evaluation metrics: Precission, Recall and F1. The positive label for this classification task is set as "Halu", for our main focus is testing the model’s ability to recognize hallucination.
As for more fine-grained hallucination-type recognition, We define a correct judgment as one where both the presence of hallucination and the specific subtype of hallucination are accurately identified. For all label types, we use micro-F1 score to quantify the performances of the three classification models.
A.9 The Prompt Designed for Detection
In Figure 13, we show the whole prompt specially designed for hallucination detection of the baselines: LLaMa-30B, Vicuna-33B, Gemini1.5 PRO, ChatGPT3.5, and GPT4. It is worth noting that due to the poor instruction-following ability and the disorderly output format of the LLaMa and Vicuna models, we conduct experiments in a 1-shot manner.
Previous works prove the LLMs’ ability to follow complex instructions (Manakul et al., 2023; Guan et al., 2023; Chen et al., 2023b; Mündler et al., 2023; Li et al., 2023b; Wang et al., 2023b; Liu et al., 2023; Li et al., 2024b), including some hallucination tasks. Thus, we reference those prompts that classify hallucination using LLMs, and then we formulate prompts for ours.
@@ A.9 The Prompt Designed for Detection
A.10 The Prompt Designed for Fine-grained Recognition
In Figure 14, we show the whole prompt manually created for fine-grained hallucination-type recognition of the three closed-source baselines: Gemini1.5 PRO, ChatGPT3.5, and GPT4.
Figure 14: The whole prompt for fine-grained hallucination-type recognition.
A.11 The Settings for CoT and Retrieval
Chain-of-Thought (CoT) (Feng et al., 2023a) describes the organized sequence of logical reasoning that unfolds during thinking. Retrieval (Gao et al., 2023; Ye et al., 2024) means retrieving relevant contents from the media to supplement external knowledge for LLMs. We employ CoT in all four domains to enhance the performance of ChatGPT3.5 and GPT4. The specific CoT is illustrated in Figure 15. Even so, only knowledge-grounded and reasoning domains are tested with retrieval via Google [^10] https://console.cloud.google.com/apis/library. This is because domains of task-oriented and chit-chat mainly involve scenarios related to daily life or virtual worlds, without specific domain knowledge as supplementary information.
@@ A.11 The Settings for CoT and Retrieval
A.12 Future Works
The necessity of dialogue-level
The sentence-level, passage-level and dialogue-level hallucination differ in the hallucination types and detection difficulties. We will explain this with the following examples to make it more clear.
Here is a sentence-level hallucination example from the dataset FactCHD (Chen et al., 2023c). ‘User: Can you tell me which mountain range is longer, the Alps or the Pyrenees? LLMs: The Pyrenees are longer than the Alps.’ In this sentence-level example, the LLM only responds to the user’s question with one sentence. We need to assess whether there are hallucination in the single sentence generated by the LLM.
Here is a passage-level hallucination example from the dataset WikiBio (Manakul et al., 2023). "Matthew Aylmer, 1st Baron Aylmer was an Irish soldier and colonial administrator. He was born in Dublin, the son of a barrister, and was educated at Trinity College, Dublin. … He was buried in Westminster Abbey." This passage-level example is directly generated by the LLM. We need to determine whether the passage with multiple sentences involves hallucination. In this example passage, hallucination occurs in the last sentence which provides unfactual information. Since there are interrelations or dependencies between the sentences in the passage, passage-level hallucination detection is more challenging than the sentence-level one.
The examples of dialogue-level hallucination are shown in Figure 2, which covers four domains and five hallucination types. The differences between it with sentence-level and passage-level hallucination are as follows: First, more types of hallucination occur frequently in dialogue. One possible reason for most benchmarks merely focusing on detecting factuality hallucination is that they are organized at a sentence or passage level. In this setting, faithfulness hallucinatios (including Incoherence, Irrelevance, and Overreliance) are less likely to occur. In contrast, since dialogue generation requires LLMs to have context coherence (Mishra et al., 2023), track the dialogue state (Heck et al., 2023), possess long-term memory capabilities (Zhong et al., 2024), and have the ability to recognize topic shifts (Lin et al., 2023), faithfulness hallucination (including Incoherence, Irrelevance, and Overreliance as described in lines 261-263) occur more frequently in dialogue. In our benchmark, the faithfulness hallucination mainly accounts for Task-oriented dialogue and Chit-Chat dialogue as shown in Figure 4. Second, it is more challenging to detect hallucination from dialogue than a single sentence or a passage. Since a dialogue contains multiple rounds of interactions that are interdependent, it can not determine whether the current round has hallucination merely based on the current content. For example, the Incoherence hallucination type occurs as the answer is not consistent with the previous context in the task-oriented dialogue in Figure 2. Thus, it needs to analyze the context dependency in the whole dialogue and judge coherence, relevance and reasoning correctness, spanning multiple rounds of interactions for dialogue-level hallucination detection, which is more challenging than sentence-level and passage-level detection.
Therefore, it is necessary to construct a dialogue-level hallucination evaluation benchmark to promote research in LLM studies.
@@ A.12 Future Works
is an important work in the future. We propose the first dedicated dialogue-level hallucination detection evaluation benchmark for LLMs and experimental results show that it is a very challenging task. Therefore, combining previous works (Chen et al., 2017; Deriu et al., 2021), developing methods based on this dataset to achieve a relatively high recognition accuracy is highly valuable.
@@
is an extension task of this work. Most existing hallucination elimination methods primarily focus on sentence-level or passage-level factuality hallucination (Luo et al., 2024). Hallucination elimination at the dialogue level not only requires models to have much parameter knowledge, but also a long-context memory capabilities, the abilities to recognize changes in topics/roles and logical transitions in the dialogue. These are helpful in addressing faithfulness hallucination. At the same time, improving the accuracy of knowledge in the LLMs’ parameters and the reasoning abilities are equally important.
@@
is the phenomenon that LLMs tend to accumulate hallucination rather than self-correcting during the generation process (Zhang et al., 2023c). Some previous works validate this phenomenon (Azaria and Mitchell, 2023; Ang et al., 2023). In our benchamark, there is a noticeable issue of hallucination snowballing. Through the experimental results, we also display such phenomenon. Eliminating hallucination snowballing in LLMs is extremely urgent in the future.
@@
During the annotation process of this dataset, we introduce a hallucination category termed "overreliance", which represents answering unanswerable content (Slobodkin et al., 2023; Sulem et al., 2021). This phenomenon signifies that LLMs tend to trust the input provided by users. Sometimes, even when there are errors in user input, the LLMs still fail to recognize them. A few past researches explore the related areas and try to find a solution. However, this issue in the application of human-machine interaction and multi-agent scenarios still remains crucial.